OpenStudy (bbb911):

2/3y -5/4y +8= -11/12y-4

OpenStudy (anonymous):

are we solving for y?

OpenStudy (bbb911):

yes solving equations

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 }y - \frac{ 5 }{ 4 }y +8 = -\frac{ 11 }{ 12}y - 4\] is this the equation?

OpenStudy (bbb911):

yes

OpenStudy (anonymous):

okay so the first thing we want to do is get all those fractions to have the same least common denominator, which is 12 so now we get: \[\frac{ 8 }{12 }y - \frac{ 15 }{ 12 }y + 8 = -\frac{ 11 }{ 12}y - 4\]

OpenStudy (bbb911):

oky

OpenStudy (anonymous):

lets get all the y's on one side and subtract/add the coefficients

OpenStudy (anonymous):

\[\frac{ 4 }{ 12 }y + 8 = -4\]

OpenStudy (anonymous):

subtract both sides by 8 \[\frac{ 4 }{12 }y = -12\]

OpenStudy (anonymous):

4/12 is actually 1/3 so \[\frac{ 1 }{ 3 }y = -12\]

OpenStudy (bbb911):

what i have is 4/12 y + -3/12y + 11/12 -12

OpenStudy (anonymous):

you changed your fractions wrong

OpenStudy (bbb911):

oh ok one sec

OpenStudy (anonymous):

2/3 ---> 8/12 5/4 ----> 15/12

OpenStudy (anonymous):

y = -36

OpenStudy (bbb911):

to get 8/12 did you multipy the 4 to 2/3?

OpenStudy (anonymous):

yes, we want that 3 to become a 12, but we have to multiply the top by 4 also so 2*4 = 8 and 3*4 = 12.... so we get 8/12

OpenStudy (bbb911):

okay what do i do with the + 8

OpenStudy (anonymous):

just move it to the other side by subtracting 8 from both sides. look at my steps above, i already solved the entire problem

OpenStudy (bbb911):

and do i subract 8 from -11/12

OpenStudy (anonymous):

no, -11/12 is a coefficient of y, so you can only subtract it from -4